OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

The slope of your line y=2x+2 is of course 2. A line perpendicular to this will have a slope of -1/2 (the negative reciprocal). Now you can use the point slope formula with the point (6,3) and slope -1/2 to get your answer: y-3=-1/2(x-6)

OpenStudy (across):

Perpendicularity of two slopes means that one is the negative reciprocal of the other. That is, a slope parallel to \(2\) is \(-1/2\).\[\]

OpenStudy (across):

Test1 is spot-on.

OpenStudy (anonymous):

Remember the point slope formula says that the equation of a line with slope m through the point (a,b) is y-b=m(x-a)

OpenStudy (anonymous):

Thank you smart people!